OpenStudy (anonymous):

QRST is a rectangle If RT=4x-16 and SQ=x+5 find the value of x

OpenStudy (anonymous):

x=7

OpenStudy (anonymous):

if u draw the figure, RT and SQ turn out to be the diagonals of the rectangle since diagonals of a rectangle are equal, so RT = SQ ==> 4x - 16 = x + 5 ==> 4x - x = 5 + 16 ==> x = 21/3 = 7

OpenStudy (anonymous):

and this time, Harkirat is correct ^_^